The perimeter of a rectangle is 230 inches. The length exceeds the width by 47 inches. Find the length and the width.
The length is _____ inches.


Thank you =)

Answer by rdamanik(3)

2L + 2W = 230 Inch
L = W + 47 Inch
L - W = 47 Inch
2L + 2W = 230 )x1 2L + 2W = 230
L - W = 47 )x2 2L - 2W = 94 +
------------------
4L = 324
L = 81
W = 81 - 47 = 34